# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

### v1

@@ VII Application of MY-based Kohn–Sham inversion to periodic settings
Beyond refinements of the inversion algorithm itself, a number of possible choices in the mathematical setup have not yet been fully explored as well. This includes a different choice than $H_{\mathrm{per}}^{-1}(\Omega)$ for the density space $X$ and $H_{\mathrm{per}}^{1}(\Omega)$ for the potential space $X^{*}$. For example, the inverse choice ($H_{\mathrm{per}}^{-1}(\Omega)$ for the potential and $H_{\mathrm{per}}^{1}(\Omega)$ for the density) would in principle work as well and exactly fits to recent results about $v$-representability on periodic one-dimensional domains [17; 18]. Although it has to be noted that this will break the self-adjointness requirement from Assumption 4 in more than one dimension. We believe that systematically exploring the question of function spaces is a promising endeavor to precisely target the inversion procedure towards reproducing the physically relevant features of the potential: as visible from Eq. (130), the function space choice is directly reflected in the norms employed during the proximal point computation. From a physical point of view, the regions close to the nucleus are generally less relevant, such that an ideal function space $X$ should feature a norm that puts less weight on the core and larger weight on the valence regions.
@@
In closing, we remark that the exceptionally close agreement between the plane-wave discretization basis and the Fourier characterization of the periodic Sobolev spaces make the periodic setting ideally suited for further joint numerical and analytical study of MY-regularized DFT. In combination with a flexible tool such as DFTK [94], mathematical ideas can be seamlessly implemented and tested both on toy problems as well as realistic materials systems. Considering specifically the similarity of MY-based KS inversion to other inversion approaches (such as the ZMP method), see Section IV, as well as the critical lack in mathematical analysis for practical inverse KS, this provides an accessible framework to close this important gap.
@@ VIII Outlook
This work aimed at showing how, since its introduction into the field of DFT, the technique of MY regularization grew ever more useful and versatile. At this point, it is interesting to ask, which are the most promising routes to follow and what more can be expected. We here collect a number of ideas and possible research directions that we deem especially interesting and worthwhile to investigate. This will at the same time serve as a summary of the presented concepts relating to MY regularization applied to DFT.
  • In Section VI it was demonstrated how in certain mean-field approximations a choice of spaces for densities and potentials that is adapted such that $\|\cdot\|_{X^{*}}^{2}$ yields the field energy automatically leads to MY regularization. A very similar choice of spaces was made in the ZMP method for density–potential inversion (Section IV) and in the numerical discussion of the periodic setting (Section VII). In general, we believe that fitting the mathematical setting to physical principles, while taking all the functional-analytic considerations of Section II.4 into account, is extremely useful as a guiding principle. The original Lieb setting [3] with $X=L^{1}\cap L^{3}$ in some sense also had this in mind via the estimate of the $L^{3}$-norm of the density by the kinetic energy, but this is just comparative, while also an exact relation to the field energy is possible. Equation (102) demonstrated how this leads to an equivalence between the duality map and the electrostatic Poisson equation. It is thus possible to encode important physical relations fully into the geometry of the density and potential spaces. (Note that such an encoding of physical laws into geometry is also present quite generally in gauge theories through the introduction of fibre bundles.) This basic idea will be pursued further for a new formulation of QEDFT in a forthcoming publication.
@@
  • A generalization in the opposite direction is also possible. This means considering Banach spaces and adapting the MY regularization itself to what is called the ‘modulus of convexity’ of the respective space. This is possible by replacing the $(\cdot)^{2}$ in Eq. (15) by $(\cdot)^{p}$, $p>1$, or right away by a Young function $\phi(\cdot)$. One then also needs to switch to a corresponding generalized duality map according to Eq. (6). While this direction was explored mathematically [26; 108; 52], it was not really implemented into the study of DFT yet.
@@
  • The original motivation for devising the regularized KS scheme was to be able to prove this convergence. Yet, clearly, a guaranteed convergence is also beneficial for practical applications, so one could try and develop functional approximations $v_{\mathrm{xc},\varepsilon}(\rho)$ that are already adapted to the regularized KS iteration of Section V.1. A small step in this direction was presented in Section V.2 that demonstrated a way how to transform available exchange–correlation potentials such that they can be applied in the regularized KS scheme. But this must be seen just as a first approach to the question. Instead a construction of such approximations from first principles could also be tried. On the other hand, the practical implications of switching to the regularized form of the KS method have not really been explored yet.
@@
  • Speaking of convergence, to date we do not have a convergence proof for the proximal-point iteration (Section IV.1) that is also the basis for the inverse KS method applied to periodic settings (Section VII). We can maybe expect it in the form of a fixed-point proof that is typical for Banach-space structured. We also note at this point, that the presented method can easily be adapted to other DFT settings, like spin DFT, current DFT, or QEDFT.
@@
  • As mentioned, the homogeneous Sobolev spaces introduced in Section VI and their duals might have a central role in future reformulations of (regularized) DFT, but they are not that well studied in functional analysis, especially not on the dual side. What is known is that they are Hilbert spaces and thus Assumption 3 is secured. We noted in Section III.3 that they are already used in some practical schemes to enhance the Kohn–Sham method. Further functional theoretic insights on these spaces would thus be a valuable contribution to ongoing work. In the course of such study, the important questions of self-adjointness and boundedness below of the Hamiltonian and, in particular, the lower semicontinuity of the ensemble constrained-search density functional must also be clarified in order to have the correct setting for MY-regularized DFT according to Assumption 4.
@@
  • The whole regularized KS framework in Section V.1 used the difference of the interacting functional $F^{1}_{\varepsilon}$ and the non-interacting $F^{0}_{\varepsilon}$, as typically done in KS-DFT. Yet, it was noted in Section VII that switching to a different guiding functional that already includes a priori information might be beneficial at least numerically. In a way we did this already in Section II.4 through the inclusion of the offset potential $v_{\mathrm{o}}$ that would include information about the external potential. But other, more complicated terms, like the Hartree mean-field energy $E_{\rm H}(\rho)$ could be included as well before regularization, thus choosing $(F^{0}+E_{\rm H})_{\varepsilon}$ as a regularized guiding functional. One only has to make sure that the functional is still convex and lower semicontinuous before regularization, in order to stay within Assumption 4. Further studies about this questions are already underway.
@@
  • In the periodic setting, the computation of proximal points, respectively the inversion of Kohn–Sham densities based on MY techniques has recently become feasible for practical systems, see Section IV. However, the currently considered algorithm still suffers from a number of restrictions: (i) it is only applicable to insulating systems, (ii) proper preconditioner have not yet been developed and (iii) the choice of regularization parameters $\varepsilon$ requires careful manual tuning. Overcoming these restrictions is crucial to broaden the scope of MY-based numerical experiments for mathematical research on DFT as well as for practical KS inversion computations.
@@
  • While recasting KS inversion via MY regularization yields an a priori orbital-free density-functional formulation, see Section IV, practical implementations of MY-based KS inversion (MYiKS) still rely on an explicit orbital parametrization to solve the proximal-point optimization problem Eq. (130). The same is true for many standard KS inversion algorithms [2; 109; 67]. Consequently, it would be advantageous to develop a mathematical framework in which the regularization is expressed directly in terms of the orbitals to provide a closer match between MY-based theory and practical inversion schemes. We remark that such an orbital-based formulation is also more naturally suited to provide a fully consistent treatment of the nonlocal term present in most pseudopotential approximations (see Section VII).
@@
  • As previously discussed in Section VII, the close match between the mathematical framework and the numerical implementation in the periodic setting is remarkable. This enables joint analytical and numerical investigations where mathematical developments can be guided and tested using numerical experiments, and suggested theoretical relationships can be directly exploited in the development of inversion algorithms with improved efficiency and reliability. Considering the many possibilities for formulating MY-regularized DFT (e.g., the considerable flexibility in selecting appropriate function spaces), we further expect numerical investigations to be decisive for suggesting which theoretical extensions are worthwhile pursuing. Moreover, we remark that since the MY-based Kohn–Sham inversion framework can be viewed as a unified framework encompassing a wide range of inversion schemes, see Section IV, we expect the resulting mathematical and algorithmic advances from such studies to be broadly applicable and to directly contribute to the development of reliable and computationally efficient inversion schemes.
@@
No new data were created or analysed in support of this research.

### v2

@@ VII Application of MY-based Kohn–Sham inversion to periodic settings
Beyond refinements of the inversion algorithm itself, a number of possible alternative choices in the mathematical setup have neither been fully explored. This includes a different choice than $H_{\mathrm{per}}^{-1}(\Omega)$ for the density space $X$ and $H_{\mathrm{per}}^{1}(\Omega)$ for the potential space $X^{*}$. For example, the inverse choice ($H_{\mathrm{per}}^{-1}(\Omega)$ for the potential and $H_{\mathrm{per}}^{1}(\Omega)$ for the density) would in principle work as well and exactly fits to recent results about $v$-representability on periodic one-dimensional domains [94; 24]. Yet, it has to be noted that this will break the self-adjointness requirement from Assumption 4 in more than one dimension [46]. We believe that systematically exploring the question of function spaces is a promising endeavor to precisely target the inversion procedure towards reproducing the physically relevant features of the potential: as visible from Eq. (131), the function space choice is directly reflected in the norms employed during the proximal point computation. From a physical point of view, the regions close to the nucleus are generally less relevant, such that an ideal function space $X$ should feature a norm that puts less weight on the core and larger weight on the valence regions.
@@
In closing, we remark that the exceptionally close agreement between the plane-wave discretization basis and the Fourier characterization of the periodic Sobolev spaces make the periodic setting ideally suited for further joint numerical and analytical study of MY-regularized DFT. In combination with a flexible tool such as DFTK [44], mathematical ideas can be seamlessly implemented and tested both on toy problems as well as realistic materials systems. Considering specifically the similarity of MY-based KS inversion to other inversion approaches (such as the ZMP method), see Section IV, as well as the critical lack of mathematical analysis for practical inverse KS, this provides an accessible framework to close this important gap.
@@ VIII Outlook
This work aimed at showing how, since its introduction into the field of DFT, the technique of MY regularization grew ever more useful and versatile. At this point, it is interesting to ask, which are the most promising routes to follow and what more can be expected. Here, we collect a number of ideas and possible research directions that we deem especially interesting and worthwhile to investigate. This will at the same time serve as a summary of the presented concepts relating to MY regularization applied to DFT.
  • In Section VI it was demonstrated how in certain mean-field approximations a choice of spaces for densities and potentials that is adapted such that $\|\cdot\|_{X^{*}}^{2}$ yields the field energy automatically leads to MY regularization. A very similar choice of spaces was made in the ZMP method for density–potential inversion (Section IV) and in the numerical discussion of the periodic setting (Section VII). In general, we believe that fitting the mathematical setting to physical principles, while taking all the functional-analytic considerations of Section II.4 into account, is extremely useful as a guiding principle. The original Lieb setting [63] with $X=L^{1}\cap L^{3}$ in some sense also had this in mind via the estimate of the $L^{3}$-norm of the density by the kinetic energy, but this is just comparative, while an exact relation to the field energy is possible. Equation (102) demonstrated how this leads to an equivalence between the duality map and the electrostatic Poisson equation. It is thus possible to encode important physical relations fully into the geometry of the density and potential spaces. (Note that such an encoding of physical laws into geometry is also present quite generally in gauge theories through the introduction of fibre bundles.) This basic idea will be pursued further for a new formulation of QEDFT in a forthcoming publication, where the adaption of the norm to the field energy means that all of Maxwell’s equations will be intimately linked to the theory. This then gives rise to the exact macroscopic Maxwell formulation of non-relativistic QED and has important consequences for renormalization.
@@
  • A generalization in the opposite direction is also possible. This means considering Banach spaces and adapting the MY regularization itself to what is called the ‘modulus of convexity’ of the respective space. This is possible by replacing the $(\cdot)^{2}$ in Eq. (15) by $(\cdot)^{p}$, $p>1$, or right away by a Young function $\phi(\cdot)$. One then also needs to switch to a corresponding generalized duality map according to Eq. (6). While this direction was explored mathematically [3; 4; 76], it was not really implemented into the study of DFT yet.
@@
  • The original motivation for devising the regularized KS scheme was to be able to prove its convergence. Yet, clearly, a guaranteed convergence is also beneficial for practical applications, so one could try and develop functional approximations $v_{\mathrm{xc},\varepsilon}(\rho)$ that are already adapted to the regularized KS iteration of Section V.1. A small step in this direction was presented in Section V.2 that demonstrated a way how to transform available exchange–correlation potentials such that they can be applied in the regularized KS scheme. But this must be seen just as a first approach to the question. Instead, a construction of such approximations from first principles could also be tried. On the other hand, the practical implications of switching to the regularized form of the KS method have not really been explored yet.
@@
  • Speaking of convergence, to date we do not have a convergence proof for the proximal-point iteration (Section IV.1) that is also the basis for the inverse KS method applied to periodic settings (Section VII). We can maybe expect it in the form of a fixed-point proof that is typical for Banach-space structures. We also note at this point that the presented method can easily be adapted to other DFT settings, like spin DFT, current DFT, or QEDFT.
@@
  • As mentioned, the homogeneous Sobolev spaces introduced in Section VI and their duals might have a central role in future reformulations of (regularized) DFT, but they are not that well studied in functional analysis, especially not on the dual side. What is known is that they are Hilbert spaces and thus Assumption 3 is secured. We noted in Section III.3 that they are already used in some practical schemes to enhance the Kohn–Sham method. Further functional theoretic insights on these spaces would thus be a valuable contribution to ongoing work. In the course of such study, the important questions of self-adjointness and boundedness below of the Hamiltonian and, in particular, the lower semicontinuity of the ensemble constrained-search density functional must also be clarified in order to have the correct setting for MY-regularized DFT according to Assumption 4. In Section II.4 it was noted that possible strategies to have a lower semicontinuous $F_{\mathrm{DM}}^{\lambda}(\rho)$ consist in choosing a bounded domain $\Omega$ (where it is possible to add periodic boundary condition) or to include an offset potential like $v_{\mathrm{off}}(\bm{r})=|\bm{r}|^{2}$. Both options are currently investigated.
@@
  • The whole regularized KS framework in Section V.1 used the difference between the interacting functional $F^{1}_{\varepsilon}$ and the non-interacting $F^{0}_{\varepsilon}$, as typically done in KS-DFT. Yet, it was noted in Section VII that switching to a different guiding functional that already includes a priori information might be beneficial at least numerically. In a way we did this already in Section II.4 through the inclusion of the offset potential $v_{\mathrm{off}}$ that would include information about the external potential. But other, more complicated terms, like the Hartree mean-field energy $E_{\rm H}(\rho)$ could be included as well before regularization, thus choosing $(F^{0}+E_{\rm H})_{\varepsilon}$ as a regularized guiding functional. One only has to make sure that the functional is still convex and lower semicontinuous before regularization, in order to stay within Assumption 4. Further studies about these questions are already underway.
@@
  • In the periodic setting, the computation of proximal points, respectively the inversion of Kohn–Sham densities based on MY techniques has recently become feasible for practical systems, see Section VII. However, the currently considered algorithm still suffers from a number of restrictions: (i) it is only applicable to insulating systems, (ii) proper preconditioners have not yet been developed, and (iii) the choice of regularization parameter $\varepsilon$ requires careful manual tuning or an explicit convergence study. Overcoming these restrictions is crucial to broaden the scope of MY-based numerical experiments for mathematical research on DFT as well as for practical KS inversion computations.
@@
  • While recasting KS inversion via MY regularization yields an a priori orbital-free density-functional formulation, see Section IV, practical implementations of MY-based KS inversion (MYiKS) still rely on an explicit orbital parametrization to solve the proximal-point optimization problem Eq. (131). The same is true for many standard KS inversion algorithms [112; 109; 91]. Consequently, it would be advantageous to develop a mathematical framework in which the regularization is expressed directly in terms of the orbitals to provide a closer match between MY-based theory and practical inversion schemes. We remark that such an orbital-based formulation is also more naturally suited to provide a fully consistent treatment of the nonlocal term present in most pseudopotential approximations (see Section VII).
@@
  • As previously discussed in Section VII, the close match between the mathematical framework and the numerical implementation in the periodic setting is remarkable. This enables joint analytical and numerical investigations, where mathematical developments can be guided and tested using numerical experiments, and suggested theoretical relationships can be directly exploited in the development of inversion algorithms with improved efficiency and reliability. Considering the many possibilities for formulating MY-regularized DFT (e.g., the considerable flexibility in selecting appropriate function spaces), we further expect numerical investigations to be decisive for suggesting which theoretical extensions are worthwhile to pursue. Moreover, we remark that since the MY-based Kohn–Sham inversion framework can be viewed as a unified framework encompassing a wide range of inversion schemes, see Section IV, we expect the resulting mathematical and algorithmic advances from such studies to be broadly applicable and to directly contribute to the development of reliable and computationally efficient inversion schemes.
@@
No new data were created or analyzed in support of this research.